# TC/TCL and NETFD Correspondence Principles for Information Backflow: A Structural Monotonicity Theorem and Minimal Phase Diagrams

### v1

@@ Abstract
We propose a unified structural framework for information backflow in minimal non-Markovian relaxation processes. The central idea is to interpret non-Markovianity as an embedding phenomenon in which observable degrees of freedom exchange information with hidden auxiliary sectors. Our approach is based on two standard formalisms in nonequilibrium physics: (i) the time-convolution (TC) and time-convolutionless (TCL) projection-operator master equations, and (ii) the correspondence principle of non-equilibrium thermo field dynamics (NETFD), which provides a doubling construction equivalent to an embedding of density-operator dynamics into a vector evolution in an enlarged space. We define a general information-backflow functional $N_{I}$ associated with an information quantity $I(t)$ and derive sufficient conditions for the absence of backflow ($N_{I}=0$) in terms of the instantaneous TCL generator. We further introduce a decomposition of backflow into classical and intrinsic (thermo-field) entanglement contributions, leading to a classification of transient overshoot and revival phenomena in a model-independent manner. Minimal classical and quantum two-state models are discussed as analytically tractable examples [19, 20], recovering Mittag–Leffler fractional relaxation as a universal envelope.
@@ 1 Introduction
  • We explicitly demonstrate how previously obtained minimal-model phase diagrams and thermo-field entanglement results are recovered as special cases of the general TC/TCL+NETFD framework [19, 20].

### v2

# Structural Theory of Information Backflow in Non-Markovian Relaxation: TC/TCL Formalism and Minimal Phase Diagrams

@@ Abstract
We develop a structural theory of information backflow in minimal non-Markovian relaxation processes within the framework of nonequilibrium statistical mechanics. The approach is based on the time-convolution (TC) and time-convolutionless (TCL) projection-operator formalisms for reduced dynamics and on the doubling construction of non-equilibrium thermo field dynamics, which provides an embedding representation of dissipative evolution.
We introduce a general backflow functional associated with a time-dependent information measure and derive generator-based sufficient conditions for the absence of backflow in terms of divisibility properties and effective relaxation rates. This allows a direct connection between memory kernels in generalized master equations and observable transient phenomena such as entropy overshoot and revival.
Furthermore, we propose a decomposition of backflow into classical mixing and intrinsic contributions in the doubled representation, leading to a unified classification of transient regimes. Minimal classical and quantum two-state models are analyzed as analytically tractable examples, yielding explicit phase diagrams and recovering Mittag-Leffler-type fractional relaxation as a universal envelope of non-Markovian damping.
The framework provides a constructive TC-to-TCL procedure for extracting effective rates and organizing memory-induced phenomena in a model-independent manner.
@@ 1 Introduction
  • We explicitly demonstrate how previously obtained minimal-model phase diagrams (Figs. 23) and thermo-field entanglement results are recovered as special cases of the general TC/TCL+NETFD framework [19, 20].
@@ 5.5 Proposition 1: Classical TCL generator criterion
Proof.
The stated conditions ensure that $\bm{W}(t)$ is a valid (time-dependent) generator of a stochastic evolution: the off-diagonal rates are nonnegative and column sums vanish, guaranteeing normalization preservation. Divisibility follows from the existence of a stochastic propagator $\bm{T}(t,s)$ satisfying $\bm{p}(t)=\bm{T}(t,s)\bm{p}(s)$ with $\bm{T}(t,s)$ stochastic for all $t\geq s$. Then the classical data-processing inequality implies $D_{\mathrm{KL}}(\bm{T}(t,s)\bm{p}\,\|\,\bm{T}(t,s)\bm{\pi})\leq D_{\mathrm{KL}}(\bm{p}\,\|\,\bm{\pi})$. If $\bm{\pi}$ is stationary, $\bm{T}(t,s)\bm{\pi}=\bm{\pi}$, which yields monotonicity of $D_{\mathrm{KL}}(\bm{p}(t)\|\bm{\pi})$. ∎
@@ 10 Discussion and Outlook
Appendix A Sharp decomposition in two-state NETFD parametrization
In the minimal two-state NETFD formulation, the extended reduced density matrix can be written in a canonical representation as
$\hat{\rho}_{A}(t)=\begin{pmatrix}p(t)&c(t)\\ c(t)^{\ast}&1-p(t)\end{pmatrix}$, (31)
where $p(t)$ denotes the population of one state and $c(t)$ encodes the intrinsic thermo-field sector.
Following the parametrization used in Ref. [19], we define the intrinsic contribution parameter
$b_{\mathrm{qe}}(t):=|c(t)|^{2}.$ (32)
The extended entropy can be decomposed as
$\hat{S}(t)=S_{\mathrm{cl}}(p(t))+S_{\mathrm{qe}}(b_{\mathrm{qe}}(t)),$ (33)
where $S_{\mathrm{cl}}$ is the classical mixing entropy and $S_{\mathrm{qe}}$ is a monotone function of $b_{\mathrm{qe}}$.
Taking the time derivative gives
$\dot{\hat{S}}(t)=S_{\mathrm{cl}}^{\prime}(p)\dot{p}(t)+S_{\mathrm{qe}}^{\prime}(b_{\mathrm{qe}})\dot{b}_{\mathrm{qe}}(t).$ (34)
In damped oscillatory regimes characteristic of fractional or memory-induced relaxation, both $p(t)$ and $b_{\mathrm{qe}}(t)$ are governed by the same relaxation envelope. Their positive-slope intervals therefore coincide, which yields the sharp additivity relation
$N_{\hat{S}}=N_{\mathrm{cl}}+N_{\mathrm{qe}}.$ (35)
This completes the argument underlying the sharp decomposition proposition.